A quite day - with no accidents
One of the attempts to improve the antenna was to try with an antenna of aluminum poles. The long aluminum, poles is placed above the surface on bamboo poles..
The science trench reached into Bølling-Allerød ice today, the drillers drilled run 500, the methan Picarro instrument is doing well and the CFA is getting into routine with the new manning. The weather has been good with high winds.
What we have done today:
1. Drilling with the NEEM long drill: 9.44 m. Drillers depth: 1470.48m
2. Logging 9.32 m of core. Logging depth 1483.94 m
3. CFA analysis: 0.55 m. Depth: 592.9 m.
4. Processing 16.5 m. Processing depth: 1449.80 m (bag 2636)
5. Improved office with storage boxes along the sides.
6. Tested HF radio and various antenna’s
Ad 1: Drillers report August 02nd : “We continue to train new drillers and to try various hallow shaft/booster configurations to learn about the drill’s performance. The closed hallow shaft configuration with no lower valve did not result in stable drilling. We abandoned this test after three short runs with high current. Afterwards we completely cleaned the hallow shaft and repaired the filter and returned the drill to the open hallow shaft configuration and reinstalled the lower valve. Full-long runs were possible again. As before, though, at least one point in each run high current demand is required, probably due to the distribution of the chips in the chip chamber. To learn why this might be happening we have repositioned the middle booster above the central junction. The result of this small change was to redistribute the chips to the upper half of the chip chamber leaving the lower third and the pump clean after a full 3.5 meter run, but still high current is needed at some point during the run that limits the motor RPMs.
Another small change was to replace the lower booster (just in front of the pump) with a centering ring. Previously, on a couple of occasions chips were packed between this booster and the pump. There has been no noticeable effect yet of replacing the booster with the ring. On at least one occasion there was also packing between the pump and the ring, but in this case the pump itself was clean.
Just before Saturday night we experienced a winch control failure upon hoisting the drill after the day’s final run. The winch simply stopped itself at 800 meters depth and the inverter turned itself off and would not restart. The backup winch control was connected without issue (Thank you Jakob and Simon). It seemed that the heating unit of the winch control had stopped working. When we tried to restart it the fan and heater did not turn on. However, Sunday morning I reconnected the original winch control and it restarted without issue. Axel and I will look into what happened more closely, but at the moment all systems are go for both winch controls".
Inclination is 2.6 degrees.”
Weather: A sunny and windy day. Blue sky, temperature between -16 to -7° C. Wind 14-18 kn from SE.
A quiet camp picture.
FL, Dorthe Dahl-Jensen
|← Previous entry||Next entry →|